# On the origin of circular rolls in rotor-stator flow

### v1

@@ Abstract
Rotor-stator flows are known to exhibit instabilities in the form of circular and spiral rolls. While the spirals are known to emanate from a supercritical Hopf bifurcation, the origin of the circular rolls is still unclear. In the present work we suggest a quantitative scenario for the circular rolls as a response of the system to external forcing. We consider two types of axisymmetric forcing: bulk forcing (based on the resolvent analysis) and boundary forcing using direct numerical simulation. Using the singular value decomposition of the resolvent operator the optimal response is shown to take the form of circular rolls. The linear gain curve shows strong amplification at non-zero frequencies following a pseudo-resonance mechanism. The optimal energy gain is found to grow rapidly with the Reynolds number (based on the rotation rate and interdisc spacing $H$) in connection with huge levels of non-normality. The results for both types of forcing are compared with former experimental works and previous numerical studies. Our findings suggest that the circular rolls observed experimentally are the combined effect of the high forcing gain and the roll-like form of the leading response of the linearised operator. For high enough Reynolds number it is possible to delineate between linear and nonlinear response. For sufficiently strong forcing amplitudes, the nonlinear response is consistent with the self-sustained states found recently for the unforced problem. The onset of such non-trivial dynamics is shown to correspond in state space to a deterministic leaky attractor, as in other subcritical wall-bounded shear flows.
@@ 1 Introduction
The spiral arms are however not the coherent structures appearing at the lowest rotation rates. Concentric rolls of finite-amplitude have been frequently reported as the earliest manifestation of unsteadiness in moderate-to-large aspect ratios. This phenomenon was first reported in spin-down experiments (Savas, 1987) and then in most experimental studies (Gauthier et al., 1999; Schouveiler et al., 2001), but have been missed by most computational stability analyses. The presence of these additional flow structures can potentially undermine all predictions from linear stability analysis. The concentric rolls were reproduced in direct numerical simulations but were reported as short transients rather than sustained coherent structures (Lopez et al., 2009). As shown by Daube & Le Quéré (2002) and recently confirmed by Gesla et al. (2023) in a strictly axisymmetric context, the base flow is linearly unstable to the axisymmetric mode for much higher Reynolds number $Re$ than suggested from experimental observation. Such an observation is consistent with the outcomes of previous linear stability studies that spirals as the first mode of linear instability. Chaotic subcritical solutions were actually identified by Daube & Le Quéré (2002). However their continuation towards lower $Re$ has failed to reach the low values relevant for circular rolls in experiments (Gesla et al., 2023). The riddle of the dynamical origin of the circular rolls reported in these experiments remains unsolved.
@@
A first hint comes from experimental study of Gauthier et al. (1999). They observed that a change of motor in their experimental set-up lowers the threshold of appearance of rolls by roughly half. This pointed in turn to a high sensitivity of the rolls to external disturbances. Following these observations, numerical computations were performed where the system was continuously forced with a sinusoidal libration of the rotor (Lopez et al., 2009; Do et al., 2010). This forcing also proved to sustain a roll-like response although the exact temporal dynamics would remain to be compared to its experimental counterpart. In particular, Do et al. (2010) demonstrated that nonlinear effects contributed to the global dynamics of the rolls. More recently it was demonstrated in Gesla et al. (2023), at least for the case $\Gamma=10$, that the axisymmetric system features, independently of any external forcing, a self-sustained nonlinear regime coexisting with the base flow below the critical threshold. The existence of this subcritical regime is reminiscent of other subcritical shear flows such as Poiseuille flows and plane Couette flow (Eckhardt, 2018; Avila & Barkley, 2023). This leads to the dilemma whether the circular rolls observed in experiments should be interpreted as a response to external forcing, in other words a noise-sustained state, or as the footprint of a self-sustained coherent state of nonlinear origin. Even in the case where the flow is externally forced, is the response linear or are nonlinear interactions important? Does the flow follow a resonance scenario or does it display its own preferred response at a given frequency, and if yes through which selection mechanism?
@@
The present study aims at answering these questions by means of resolvent analysis and numerical simulations. Although the circular rolls were reported experimentally only at low values of $Re$ where the flow stays axisymmetric, we propose to consider $Re$ as the main governing parameter for the purely axisymmetric rotor-stator configuration, regardless of whether the experimental dynamics would stay axisymmetric or not, and to investigate the effect of varying $Re$ on both the linear receptivity scenario and the fully nonlinear one. Our receptivity approach to external forcing follows several complementary approaches. Classical optimal linear response theory, through resolvent analysis, directly yields the optimal forcing eliciting the strongest response of the flow. By design however, this forcing is applied within the bulk of the flow, away from the solid walls. Parasite vibrations of the set-up are rather expected to act at the fluid-solid interface and, thus, bulk-based optimisation might not capture them especially when the forcing is modelled as an additive force. For this reason, as well as for the freedom of incorporating nonlinear fluid interactions, direct numerical simulation of fluid flow in the presence of well controlled boundary oscillations is also considered without any optimisation. As we shall see, the scenario most consistent with experimental observations is the boundary forcing. For higher $Re$, we demonstrate numerically how circular rolls of finite amplitude can be elicited and sustained by nonlinear forcing, whereas depending on the value of $Re$ they correspond either to super-transients or to an attracting dynamics, should the forcing be turned off. Beyond the immediate analogy with subcritical shear flows, it is also of strong general interest as it generalises the phenomenon of linear receptivity to the nonlinear framework.
@@ 2 Governing equations and numerical methods
The system consists of two coaxial disks of radius $R$, separated by a gap $H$. One of the disks, the rotor, rotates at a constant dimensional angular velocity $\Omega$, while the stator is at rest, see figure 1. We chose to non-dimensionalise all lengths by the gap $H$ and time using $(\Omega)^{-1}$. Assuming a constant kinematic viscosity $\nu$ for the fluid, two non-dimensional parameters characterise this system, namely the geometric aspect ratio $\Gamma=R/H$ and the (gap-based) Reynolds number $Re=\Omega H^{2}/\nu$.
@@ 3 Base flow
For the case $\Gamma=10$, a linear instability of the base flow in the axisymmetric configuration was found for $Re\approx 3000$, due to the destabilisation of a wavepacket of counter-rotating circular rolls in the Bödewadt layer at $r\approx 2$ for $Re\approx 3000$ (Daube & Le Quéré, 2002). This result was recently confirmed quantitatively by Gesla et al. (2023). This threshold for axisymmetric instability is one order of magnitude larger than the experimental threshold of $Re\approx 200$ reported e.g. in Gauthier (1998).
The base flow being linearly stable in the $Re$-interval of interest for this study, it can in principle be found by asymptotic time integration. Nevertheless, for accuracy purposes and in order to avoid to avoid strong transient growth effects (Daube & Le Quéré, 2002; Gesla et al., 2023) it is safest to converge the base flow using the dedicated Newton-Raphson solver. In the rest of the paper, once the base flow is known, we deal exclusively with axisymmetric perturbation ${\bm{u}}(r,z,t)$ to the base flow solutions, i.e. ${\bm{u}}={\bm{v}}-{\bm{U}_{b}}$ and $p=\pi-p_{b}$, the equations for which can be deduced directly from the governing equations by subtraction. The associated boundary conditions are homogeneous of Dirichlet type, i.e. ${\bm{u}}=0$ at all walls. At the axis ($r=0$) $u_{r}$ and $u_{\theta}$ vanish while $u_{z}$ verifies a Neumann condition.
@@ 4.1 Optimal response theory
The coupled system in (9)-(10) is linear in ${\bm{u}}$, $p$ and also ${\bm{f}}$, which suggests to use the resolvent formalism. The theory by Trefethen & Embree (2005) forms the ideal framework except that the original linear system needs be rewritten in a form $\partial{\bm{z}}/\partial t={\bm{L}\bm{z}}+{\bm{f}_{z}}(t)$, where the unknown field ${\bm{z}}$, a linear operator ${\bm{L}}$ and a forcing ${\bm{f}_{z}}(t)$ need to be specified. We introduce the new variable $\mathbf{q}=(u_{r},u_{\theta},u_{z},p)$ which contains the values of the fields $u_{r}$,$u_{\theta}$, $u_{z}$ and $p$ at all the points of the discretised domain. The size of $\mathbf{q}$ is $O(4N_{r}N_{z})$. We also introduce the rectangular prolongation operator ${\bm{P}}$ of size $O(4N_{r}N_{z}\times 3N_{r}N_{z})$ which maps ${\bm{u}}$ into ${\bm{q}}$, so that $\mathbf{P}^{T}\mathbf{q}=\mathbf{u}$ with the property $\mathbf{P}^{T}\mathbf{P}=\mathbf{I}$, see e.g. Jin et al. (2021). The linear system (910) can then be rewritten into the new form
@@ 4.2 Optimal response : results
Figure 3: Optimal response for bulk-based forcing. Left: optimal energy gain as a function of forcing angular frequency $\omega$, given by the largest singular value of the resolvent operator (14). Right: optimal gain across all $\omega$’s as a function of $Re$. Numerical resolution R0.
@@
For $Re\approx 150$ the optimally forced structures change radically as unsteady ($\omega\neq 0$) forcing takes over steady forcing. This is clearly seen in figure 3(right) where for $Re\gtrsim 150$ nonzero forcing frequencies $\omega$ start to dominate the optimal gain curve. The structure of the associated optimal forcing lies entirely within the Bödewadt layer, see figure 5 which focuses on $\omega\approx 1.7$ close to the optimal angular frequency (see Table 2). For all velocity components these structures are tilted by the mean shear into the streamwise direction. While the amplitude of optimal forcing is similar in the three spatial components the corresponding optimal response may differ from component to component (Jovanović & Bamieh, 2005), and in the present geometry it is clearly dominated by the azimuthal component. The signature of the circular rolls can be also seen in the azimuthal response although the corresponding structures should more realistically be labelled as azimuthal streaks. Their position ($r\approx 4-6$) and the corresponding $Re$ are perfectly consistent with experimental observations. Most importantly, and this constitutes one of the main findings of the current work, the optimal response is in the shape of circular rolls in the Bödewadt layer.
@@
The evolution of the optimally forced structure is now described as $Re$ increases beyond 250. It is first noted that, as shown in figure 3 (right), the optimal gain $G_{opt}(Re)$ grows exponentially with $Re$. We note that the experimental study of Gauthier et al. (1999), performed for $\Gamma\approx 20$, has suggested a supercritical bifurcation as the origin of the observed rolls. Our observations do not corroborate this hypothesis since the angular frequency content of the response features that a continuum of frequencies can be excited externally, at odds with the dominant frequency stemming from a Hopf bifurcation. Moreover, the steep exponential increase of $G(Re)$ reported above might be responsible, in experimental conditions where external forcing stays uncontrolled, for the apparent bifurcating behaviour where the amplitude of the response increases rapidly with $Re$. As $Re$ continues to increase, as shown in figure 6 the optimal forcing evolves with $Re$ from a wide support within the Bödewadt layer to thinner structures in both the Bödewadt and the shrouding wall, and even for $Re=3000$ also in the Ekman layer. For such high $Re$ values the respective supports of the forcing and the response are almost disjoint.
@@
The results above unambiguously point towards an unsteady response in the shape of circular rolls for all $Re\gtrsim 150$. We emphasize that the crossing of the gain curves for $\omega=0$ and $\neq 0$ in figure 3(right) does not define a threshold value for $Re$ because, as for any linear receptivity mechanism, the response depends linearly on the spectral content of the forcing history. Defining a threshold value for $Re$ is demanding because it is highly dependent on the amplitude levels that an experimentalist can detect in practice. By focusing on the value of the optimal gain ($G=4.9\times 10^{2}$) at $Re=250$ the following hand-wavy evaluation can be made. Any parasite vibration present in the experiment projected on the orthogonal basis of the optimal modes will have a nonzero component that will be optimally forced. If the amplitude of this component is, say, of order $O(10^{-2})$ it will be amplified by the linear mechanism by $\sqrt{G}\approx 20$ to yield an $O(10^{-1})$ response, which can be detected in experiments.

### v2

@@ Abstract
Rotor-stator flows are known to exhibit instabilities in the form of circular and spiral rolls. While the spirals are known to emanate from a supercritical Hopf bifurcation, the origin of the circular rolls is still unclear. In the present work we suggest a quantitative scenario for the circular rolls as a response of the system to external forcing. We consider two types of axisymmetric forcing: bulk forcing (based on the resolvent analysis) and boundary forcing using direct numerical simulation. Using the singular value decomposition of the resolvent operator the optimal response is shown to take the form of circular rolls. The linear gain curve shows strong amplification at non-zero frequencies following a pseudo-resonance mechanism. The optimal energy gain is found to scale exponentially with the Reynolds number $Re$ (for $Re$ based on the rotation rate and interdisc spacing $H$). The results for both types of forcing are compared with former experimental works and previous numerical studies. Our findings suggest that the circular rolls observed experimentally are the effect of the high forcing gain together with the roll-like form of the leading response of the linearised operator. For high enough Reynolds number it is possible to delineate between linear and nonlinear response. For sufficiently strong forcing amplitudes, the nonlinear response is consistent with the self-sustained states found recently for the unforced problem. The onset of such non-trivial dynamics is shown to correspond in state space to a deterministic leaky attractor, as in other subcritical wall-bounded shear flows.
@@ 1 Introduction
The spiral arms are however not the coherent structures appearing at the lowest rotation rates. Concentric rolls of finite-amplitude have been frequently reported as the earliest manifestation of unsteadiness in moderate-to-large aspect ratios. This phenomenon was first reported in spin-down experiments (Savas, 1987) and then in most experimental studies (Gauthier et al., 1999; Schouveiler et al., 2001), but have been missed by most computational stability analyses. The presence of these additional flow structures can potentially undermine all predictions from linear stability analysis. The concentric rolls were first reproduced in direct numerical simulations following impulsive perturbations, but they were reported to convect towards the center and to vanish, being hence short transients rather than sustained coherent structures (Lopez et al., 2009). This is consistent so far with the convective instability viewpoint put forward in the recent review by Martinand et al. (2023), although such a viewpoint is intrinsically limited in finite geometries. As shown by Daube & Le Quéré (2002) and recently confirmed by Gesla et al. (2024) in a strictly axisymmetric context, the base flow is linearly unstable to the axisymmetric mode for much higher Reynolds number $Re$ than suggested from experimental observation. This agrees with previous linear stability studies that predict spirals as the first mode of linear instability. Chaotic subcritical solutions were actually identified by Daube & Le Quéré (2002). However their continuation towards lower $Re$ has failed to reach the low values relevant for circular rolls in experiments (Gesla et al., 2024). The riddle of the dynamical origin of the circular rolls reported in these experiments thus remains unsolved.
Continuous efforts have been made in the past to link the stability of enclosed rotor-stator flows to the results of local stability analysis of one-dimensional similarity solutions. All these studies rely on boundary layer profiles and imply unbounded domains in the radial direction for one- or two-disc set-ups. In the latter case, large enough Reynolds numbers are also necessary so that the boundary layers are well separated. A recent review of transition scenarios in finite- and infinite-radius rotor-stator configurations based on a local stability analysis can be found in Martinand et al. (2023). The local stability analysis is necessarily approximative for enclosed rotor-stator flow but nevertheless very useful to understand the physical mechanism of instabilities. The Reynolds number range where the circular rolls first appear does however not correspond to separated boundary layers, which questions the quantitative relevance of local stability analysis.
@@
A first hint about the origin of the circular rolls comes from the experimental study by Gauthier et al. (1999). They observed that a change of motor in their experimental set-up lowers the threshold of appearance of rolls by roughly half. This pointed in turn to a high sensitivity of the rolls to external disturbances. Following these observations, numerical computations were performed where the system was continuously forced with a sinusoidal libration of the rotor (Lopez et al., 2009; Do et al., 2010). This forcing also proved to sustain a roll-like response although the exact temporal dynamics would remain to be compared to its experimental counterpart. Based on direct numerical simulation, Do et al. (2010) demonstrated that nonlinear effects contributed to the global dynamics of the rolls. We present results of the optimal forcing from the resolvent analysis that complement the direct numerical simulations and bring new insigths. Since the form of the forcing in Do et al. (2010) is fixed, their study can be understood as optimisation in the forcing frequency only. The form of the optimal response resulting from the resolvent analysis can be understood as a spatial optimisation and form a novelty with respect to Do et al. (2010). More recently it was demonstrated in Gesla et al. (2024), at least for the case $\Gamma=10$, that the axisymmetric system features, independently of any external forcing, a self-sustained nonlinear regime coexisting with the base flow below the critical threshold. The existence of this subcritical regime is reminiscent of other subcritical shear flows such as Poiseuille flows and plane Couette flow (Eckhardt, 2018; Avila et al., 2023). This leads to the dilemma whether the circular rolls observed in experiments should be interpreted as a response to external forcing, in other words a noise-sustained state, or as the footprint of a self-sustained coherent state of nonlinear origin. Formulated differently, does the flow follow a resonance scenario or does it oscillate autonomously independently of the way it is forced? Even in the case where the flow is externally forced, is the response linear or are nonlinear interactions important?
@@
The present study aims at answering these questions by means of resolvent analysis and numerical simulations. Although the circular rolls were reported experimentally only at low values of $Re$ where the flow stays axisymmetric, we propose to consider $Re$ as the main governing parameter for the purely axisymmetric rotor-stator configuration, regardless of whether the experimental dynamics would stay axisymmetric or not, and to investigate the effect of varying $Re$ on both the linear receptivity scenario and the fully nonlinear one. Due to the assumed axisymmetric configuration the present study does not bring any insight into optimally forced 3D structures. Our receptivity approach to external forcing follows several complementary approaches. Classical optimal linear response theory, through resolvent analysis, directly yields the optimal forcing eliciting the strongest response of the flow. By design however, this forcing is applied within the bulk of the flow, away from the solid walls. Parasite vibrations of the set-up are rather expected to act at the fluid-solid interface and, thus, bulk-based optimisation might not capture them especially when the forcing is modelled as an additive force. For this reason, as well as for the freedom of incorporating nonlinear fluid interactions, direct numerical simulation of fluid flow in the presence of well controlled boundary oscillations is also considered without any optimisation. As we shall see, the scenario most consistent with experimental observations is the boundary forcing. For higher $Re$, we demonstrate numerically how circular rolls of finite amplitude can be elicited and sustained by nonlinear forcing, whereas depending on the value of $Re$ they correspond either to super-transients or to an attracting dynamics, should the forcing be turned off. Beyond the immediate analogy with subcritical shear flows, it is also of strong general interest as it yields a concrete example of nonlinear receptivity whereas receptivity is traditionally investigated using linear tools.
@@ 2 Governing equations and numerical methods
The system consists of two coaxial disks of radius $R$, separated by a gap $H$. One of the disks, the rotor, rotates at a constant dimensional angular velocity $\Omega$, while the stator is at rest, see figure 1. We chose to non-dimensionalise all lengths by the gap $H$ and time using $(\Omega)^{-1}$. Assuming a constant kinematic viscosity $\nu$ for the fluid, two non-dimensional parameters characterise this system, namely the geometric aspect ratio $\Gamma=R/H$ and the (gap-based) Reynolds number $Re=\Omega H^{2}/\nu$. Other possible definitions for the Reynolds number $Re$ include $Re_{R}=\Omega R^{2}/\nu$, built on the length scale $R$ or $Re_{RH}={\Omega RH}/{\nu}$. In the current work the definition of $Re$ based on $H$ is preferred since it can be defined in both finite and infinite radial extent set-ups, as it facilitates a straightforward comparison with other studies.
@@ 3 Base flow
For the case $\Gamma=10$, a linear instability of the base flow in the axisymmetric configuration was found for $Re\approx 3000$, due to the destabilisation of a wavepacket of counter-rotating circular rolls in the Bödewadt layer at $r\approx 2$ (Daube & Le Quéré, 2002). This result was recently confirmed quantitatively by Gesla et al. (2024). Threshold for axisymmetric instability is one order of magnitude larger than the experimental threshold of $Re\approx 200$ reported e.g. in Gauthier (1998).
The analysis of Gesla et al. (2024) can be extended to the non-axisymmetric case by using an exponential ansatz with an integer wavenumber $m$ in the azimuthal direction (Gelfgat, 2015). The instability investigation of Gesla et al. (2024) becomes spatially two-dimensional with an additional parameter $m$ (azimuthal wavenumber) and a critical value of $Re=528.91$ for $m=32$ (numerical resolution $N_{r}\times N_{z}=600\times 128$ for $\Gamma=10$ set-up with rotating shroud). It is noted that the critical $Re$ for non-axisymmetric modes is around six times lower than for axisymmetric modes ($Re\approx 3000$). The results for $m\neq 0$ concern spiral rolls rather than circular roll. Since this work focuses only on circular rolls, it is assumed that the flow is strictly axisymmetric in what follows.
The base flow being linearly stable in the $Re$-interval of interest for this study, it can in principle be found by asymptotic time integration. Nevertheless, in order to avoid strong transient growth effects (Daube & Le Quéré, 2002; Gesla et al., 2024) it is safest to converge the base flow using the dedicated Newton-Raphson solver to machine precision (i.e. when the Euclidean norm of the residual drops below $10^{-11}$). As shown in section 6, for high enough $Re$ even a small perturbation will evolve to a top branch chaotic solution. Such a perturbation occurs for instance following a instantaneous change in $Re$. It was in particular found that the base flow cannot be found in a reasonable time using time integration for $Re>2100$.
In the rest of the paper, once the base flow is known, we deal exclusively with axisymmetric perturbation ${\bm{u}}(r,z,t)$ to the base flow, i.e. ${\bm{u}}={\bm{v}}-{\bm{U}_{b}}$ and $p=\pi-p_{b}$, the equations for which can be deduced directly from the governing equations by subtraction. The associated boundary conditions are homogeneous of Dirichlet type, i.e. ${\bm{u}}=0$ at all walls. At the axis ($r=0$) $u_{r}$ and $u_{\theta}$ vanish while $u_{z}$ verifies a Neumann condition.
@@ 4.1 Optimal response theory
The coupled system in (9)-(10) is linear in ${\bm{u}}$, $p$ and also ${\bm{f}}$, which suggests to use the resolvent formalism. The theory by Trefethen & Embree (2005) forms the ideal framework except that the original linear system needs be rewritten in a form $\partial{\bm{q}}/\partial t={\bm{L}\bm{q}}+{\bm{f}_{q}}(t)$, where the unknown field ${\bm{q}}$, a linear operator ${\bm{L}}$ and a forcing ${\bm{f}_{q}}(t)$ need to be specified. We introduce the new variable $\mathbf{q}=(u_{r},u_{\theta},u_{z},p)$ which contains the values of the fields $u_{r}$,$u_{\theta}$, $u_{z}$ and $p$ at all the points of the discretised domain. The size of $\mathbf{q}$ is $O(4N_{r}N_{z})$. We also introduce the rectangular prolongation operator ${\bm{P}}$ of size $O(4N_{r}N_{z}\times 3N_{r}N_{z})$ which maps ${\bm{u}}$ into ${\bm{q}}$, so that $\mathbf{P}^{T}\mathbf{q}=\mathbf{u}$ with the property $\mathbf{P}^{T}\mathbf{P}=\mathbf{I}$, see e.g. Jin et al. (2021). The linear system (910) can then be rewritten into the new form
@@ 4.2 Optimal response : results
Figure 3: Optimal response for bulk-based forcing. Left: optimal energy gain as a function of forcing angular frequency $\omega$, given by the largest singular value of the resolvent operator (14). Right: optimal gain across all $\omega$’s as a function of $Re$. Numerical resolution R0. The inset shows the optimal forcing angular frequency as listed in table 2.
@@
For $Re\approx 150$ the optimally forced structures change radically as unsteady ($\omega\neq 0$) forcing takes over steady forcing. This is clearly seen in figure 3(right) where for $Re\gtrsim 150$ nonzero forcing frequencies $\omega$ start to dominate the optimal gain curve. The structure of the associated optimal forcing lies entirely within the Bödewadt layer, see figure 5 which focuses on $\omega\approx 1.7$ close to the optimal angular frequency (see Table 2). For all velocity components these structures are tilted by the mean shear into the streamwise direction. While the amplitude of optimal forcing is similar in the three spatial components the corresponding optimal response may differ from component to component (Jovanović & Bamieh, 2005), and in the present geometry it is clearly dominated by the azimuthal component. The signature of the circular rolls can be also seen in the azimuthal response although the corresponding structures should more realistically be labelled as azimuthal streaks. Their position ($r\approx 4-6$) is perfectly consistent with the experimental observations by Schouveiler et al. (2001) and the numerical studies of Lopez et al. (2009) and Do et al. (2010). Most importantly, and this constitutes one of the main findings of the current work, the optimal response is in the shape of circular rolls in the Bödewadt layer.
@@
The evolution of the optimally forced structure is now described as $Re$ increases beyond 250. It is first noted that, as shown in figure 3 (right), the optimal gain $G_{opt}(Re)$ grows exponentially with $Re$. Other examples of an exponential scaling of gain include a backward-facing step (Blackburn et al., 2008) and an oscillatory pipe flow (Xu et al., 2021). We note that the experimental study of Gauthier et al. (1999), performed for $\Gamma\approx 20$, has suggested a supercritical bifurcation as the origin of the observed rolls. Our observations do not corroborate this hypothesis since the angular frequency content of the response features that a continuum of frequencies can be excited externally, at odds with the dominant frequency stemming from a Hopf bifurcation. Moreover, the steep exponential increase of $G(Re)$ reported above might be responsible, in experimental conditions where external forcing stays uncontrolled, for the apparent bifurcating behaviour where the amplitude of the response increases rapidly with $Re$. As $Re$ continues to increase, as shown in figure 6 the optimal forcing evolves with $Re$ from a wide support within the Bödewadt layer to thinner structures in both the Bödewadt and the shrouding wall, and even for $Re=3000$ also in the Ekman layer. For such high $Re$ values the respective supports of the forcing and the response are almost disjoint.
@@
The results above unambiguously point towards an unsteady response in the shape of circular rolls for all $Re\gtrsim 150$. We emphasize that the crossing of the gain curves for $\omega=0$ and $\neq 0$ in figure 3(right) does not define a threshold value for $Re$ because, as for any linear receptivity mechanism, the response depends linearly on the spectral content of the forcing history. Defining a threshold value for $Re$ is demanding because it is highly dependent on the amplitude levels that an experimentalist can detect in practice. By focusing on the value of the optimal gain ($G=4.9\times 10^{2}$) at $Re=250$ the following evaluation can be made. Any parasitic vibration present in the experiment projected on the orthogonal basis of the optimal modes will have a nonzero component that will be optimally forced. If the amplitude of this component is, say, of order $O(10^{-2})$ it will be amplified by the linear mechanism by $\sqrt{G}\approx 20$ to yield an $O(10^{-1})$ response, which can be detected in experiments.